# Learned Free-Energy Functionals from Pair-Correlation Matching for Dynamical Density Functional Theory

### v1

@@ Abstract
Figure 4: Breakthrough simulations using neural DDFT and Browninan dynamics (see SM [23] for the movie version). (a) LJ particles from the source move towards the sink under two different MOF-like repulsive external potentials. The dynamics of the density evolution and the final evolved density (visualized) obtained from DDFT are in close agreement with BD. (b) Time evolution of the output to input flux ratio, showing the approach towards steady-state or breakthrough condition. The particles reach breakthrough earlier under the MOF-2 external potential. The curves obtained from DDFT are in close agreement with BD, while being smoother and taking only $2\%$ of the compute time of BD.
@@
We simulate this system using our open DDFT equation (eq. 2) for a duration of $150\tau$, with a timestep of $10^{-3}\tau$ and grid spacing of $\frac{\sigma}{32}$. To obtain a breakthrough curve we compute the time-dependence of the ratio $J(z_{2},t)/J(z_{1},T)$ of the particle fluxes $J(z_{2},t)$ close to the output at $z_{2}=7.5\sigma$ and $J(z_{1},T)$ close to the input at $z_{1}=2.5\sigma$ at late time $T=145\tau$ where steady-state is (almost) reached. The particle flux from DDFT is computed as $J(z,t)=-\Gamma\rho(z,t)\nabla\frac{\delta\mathcal{F}[\rho(z)]}{\delta\rho(z)}$, for which we also apply a Savitsky-Golay filter [30] over the spatial gradient to smooth out noise arising from numerical errors. We compare this with a BD simulation where the particle baths are modelled using hybrid Grand-Canonical Monte Carlo moves (see SM [23], Sec.3). The BD flux is computed as the net number of particles crossing the planes at $z_{1}$ and $z_{2}$ per unit time per unit area, averaged over $500$ BD runs. As observed in Fig. 4, the dynamics of both the BD and DDFT simulations agree closely and reach breakthrough while the DDFT simulation takes only $2\%$ of the compute time of the BD simulation (see SM [23], Sec.5 for runtime comparisions). The potential from MOF-2 (red), which corresponds to a MOF with larger pore size, presents a lower particle barrier which gives rise to a faster evolution to the steady state. Note that since the GCMC moves are run in between each BD update while the open DDFT equation updates are simultaneous, we do not expect exact dynamical agreement though this can be alleviated to some extent with the choice of $\Gamma^{\prime}$ in the growth term.
@@
Conclusion—To summarize, we have presented results that the neural excess free-energy functional trained via pair-correlation matching can be directly applied to DDFT without retraining, with high accuracy relative to an analytical free-energy approximation and with high efficiency relative to BD simulation for conserved dynamics, as well as for non-conserved dynamics via a novel gradient flow based extension inspired by gas adsorption simulations. Future work could focus on the extension of this method to systems with full $3$D geometry and anisotropic particles, and extending it to incorporate superadiabatic forces [31], towards practical applications in modelling non-equilibrium systems.
@@
Acknowledgements—We would like to thank David Dubbeldam for useful discussions.
@@
Data availability—Code and data to reproduce all results will be made available upon publication.

### v2

@@ VI Experiments
Figure 4: Breakthrough simulations using neural DDFT and Brownian dynamics (see Appendix E for the movie version). (a) LJ particles from the source move towards the sink under two different MOF-like repulsive external potentials. The dynamics of the density evolution and the final evolved density (visualized) obtained from DDFT are in close agreement with BD. (b) Time evolution of the output to input flux ratio, showing the approach towards steady-state or breakthrough condition. The particles reach breakthrough earlier under the MOF-2 external potential. The curves obtained from DDFT are in close agreement with BD, while being smoother and taking only $2\%$ of the compute time of BD.
We run two sets of “experiments” to evaluate the accuracy and efficiency of neural DDFT from pair-correlation matching with respect to BD, one on a closed system with planar geometry and conserved number of particles under various external potentials, and the other on an open system with a source and sink with varying number of particles, inspired by a gas adsorption process.
VI.1 Closed system
We use a 3D system of LJ particles with planar geometry, under external potentials that are mixtures of repulsive Gaussian and steep well potentials as in [10]. In addition to the neural functional from pair-correlation matching (denoted by $\mathcal{F}_{\theta,\text{exc}}^{(2)}$), we also train a functional using single-body direct correlation matching (denoted by $\mathcal{F}_{\theta,\text{exc}}^{(1)}$) (see Section A.1). Both functionals use the same convolutional neural network with periodic and dilated convolutions, each with a kernel size of 3, a dilation of 2, and 6 layers. The layer dimensions are set to $N_{\text{channels}}$ = [16, 16, 32, 32, 64, 64], with average pooling with kernel size 2 applied after each layer. The models are trained for 5000 epochs, each taking less than an hour on an RTX 4070 GPU for convergence.
For training $\mathcal{F}_{\theta,\text{exc}}^{(2)}$, we simulated $1000$ uniform densities and measured their radial distribution functions with ${\sigma}/{100}$ grid-size and chemical potentials in the range -4 to 0.5. $\mathcal{F}_{\theta,\text{exc}}^{(1)}$ is trained on $800$ non-uniform densities [10] with ${\sigma}/{100}$ grid-size and chemical potentials in the range -4 to 0.5. The dataset of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ is larger, since no validation set had to be excluded from the train set, since it is validated on samples from the non-uniform dataset. We also evaluate against an analytical functional $\mathcal{F}^{\text{MF}}_{\text{exc}}$ which models the attractive contribution using mean-field approximation and the hard-sphere contribution using White-Bear II version of FMT [33], as implemented in PyDFTlj [1].
Our experiments show that the neural free energy functionals $\mathcal{F}_{\theta,\text{exc}}^{(1)}$ and $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ are both significantly better than the $\mathcal{F}^{\text{MF}}_{\text{exc}}$ functional in accuracy across time, as shown in. Fig. 2 & Fig. 3. The convergence of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ is similar to $\mathcal{F}_{\theta,\text{exc}}^{(1)}$, except in Fig. 3(d) where the density near the wells is pushed above the maximum bulk value of $0.67\sigma^{-3}$ seen during training of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$. The errors for all the DDFT approaches for time intervals $t<1\tau$ are high which we attribute to overly fast dynamics arising from the adiabatic approximation [22]. More notably, the dynamical density profiles from DDFT are several orders of magnitude faster to compute with DDFT than with BD simulation: while running $5000$ trials of BD takes about a day on a single-core CPU, the DDFT approach on the same computer takes less than 6 minutes (see Appendix D for runtime comparisions).
VI.2 Open system
To evaluate the pair-correlation matching functional $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ on an open system, we construct a system modeled after “breakthrough” experiments used in gas adsorption studies [34, 35]. In these studies, typically a column filled with adsorbent material (e.g. nanoporous metal-organic framework (MOF) is exposed to a gas stream (e.g. CO2), where the breakthrough of the gas is detected at the outlet to estimate the adsorption capacity and kinetics of the material. We adopt such a setup by extending the closed Lennard-Jones system from the previous experiment with two particle baths, one at a high $\mu=0.5$ in $z\in[0,2]\sigma$ to simulate an LJ particle source and the other at extremely low $\mu=-100$ in $z\in[8,10]\sigma$ to simulate an LJ particle sink. The difference in $\mu$ causes particles to flow from the source to the sink through the region $z\in[2,8]\sigma$, in which we apply, in addition, a planar potential to mimic two MOF planes from the IRMOF series [36] (see Appendix C). In the reference BD simulation, we extend the closed system to an open system by attaching two Grand-Canonical Monte-Carlo (GCMC) reservoirs as source and sink. The source reservoir is from $z=[0,2]\sigma$ with $\mu=0.5$ and the sink is from $z=[8,10]\sigma$ with $\mu=-100$. 10 trials of GCMC insertion/deletion steps, with no translation/rotation, are attempted in LAMMPS (hybrid) between every BD update. Before starting the breakthrough simulation, we first randomly initialize LJ particles in the source region and confine them using a potential barrier and equilibrate the system for $20000$ steps (each timestep is $10^{-3}\tau$). Then the barrier is removed and the MOF-like potential is applied on the system and $100000$ collection steps are performed as the particles flow towards the sink. The planar density is sampled every 100 steps, over a bin size of ${\sigma}/{32}$ and is averaged over $500$ BD runs.
@@ VI.2 Open system
We simulate this same system using our open DDFT equation (eq. 8) for a duration of $150\tau$, with a timestep of $10^{-3}\tau$ and grid spacing of ${\sigma}/{32}$. To obtain breakthrough curves we compute the time-dependence of the ratio $J(z_{2},t)/J(z_{1},T)$ of the particle fluxes $J(z_{2},t)$ close to the output at $z_{2}=7.5\sigma$ and $J(z_{1},T)$ close to the input at $z_{1}=2.5\sigma$ at late time $T=145\tau$ where steady-state is (almost) reached. The particle flux from DDFT is computed as $J(z,t)=-\Gamma\rho(z,t)\nabla\left({\delta\mathcal{F}[\rho(z)]}/{\delta\rho(z)}\right)$, where $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ is used for the excess component. We also apply a Savitsky-Golay filter [37] over the spatial gradient in the flux computation to smooth out any noise arising from numerical errors. The BD flux is computed as the net number of particles crossing the planes at $z_{1}$ and $z_{2}$ per unit time per unit area, averaged over $500$ BD runs.
As observed from the breakthrough curves in Fig. 4, the dynamics of both the BD and DDFT simulations agree closely and reach breakthrough at similar times. The potential from MOF-2 (red), which corresponds to a MOF with larger pore size, presents a lower particle barrier giving rise to a faster evolution to the steady state, which is captured by both simulations. But the DDFT simulation is much more efficient, taking only $2\%$ of the compute time of the BD simulation (see Appendix D for runtime comparisons). Note that since the GCMC moves are run in between each BD update while the open DDFT equation updates are simultaneous, we do not expect exact dynamical agreement though this can be alleviated to some extent with the choice of $\Gamma^{\prime}$ in the growth term.
VII Conclusion
@@
To summarize, we have presented results that the neural excess free-energy functional trained via pair-correlation matching can be directly applied to DDFT without retraining, with high accuracy relative to an analytical free-energy approximation and with high efficiency relative to BD simulations. We show this for both conserved dynamics, as well as for non-conserved dynamics via a novel gradient flow based extension which we evaluate on a system inspired by breakthrough adsorption simulations. Future work could focus on the extension of this method to systems with full $3$D geometry and anisotropic particles, and extending it to incorporate superadiabatic forces [38], towards practical applications in modeling many-body non-equilibrium systems.
Acknowledgements
@@
We would like to thank David Dubbeldam and Christian Koke for useful discussions.
Data availability
@@
Code and data to reproduce all results will be made available upon publication.
Appendix A Classical Density Functional Theory
Classical density functional theory (cDFT) [3] proves the existence of a grand potential functional $\Omega[\rho]$ of a one-body density profile $\rho(\mathbf{r})$, with the condition that the equilibrium density profile $\rho_{\text{eq}}(\mathbf{r})$ minimizes $\Omega[\rho]$ and is equal to the equilibrium grand potential $\Omega_{0}$. From $\Omega_{0}$, all the thermodynamic properties of the system of interest can be subsequently derived.
$\left.\frac{\partial\Omega(\mu,[\rho],T)}{\partial\rho(\mathbf{r})}\right|_{\rho=\rho_{\text{eq}}}=0,\quad\Omega([\rho_{\text{eq}}])=\Omega_{0}.$ (14)
The grand potential functional $\Omega[\rho]$ can be written as,
$\Omega(\mu,[\rho],T)=\mathcal{F}([\rho],T)-\int\mathrm{d}^{3}r\rho(\mathbf{r})(\mu-V_{\text{ext}}(\mathbf{r})),$ (15)
where $\mu$ is the chemical potential, $T$ is the temperature, and $\mathcal{F}[\rho]$ is the intrinsic Helmholtz free-energy functional which can be split into three parts,
$\mathcal{F}([\rho],T)=\mathcal{F}_{\text{id}}([\rho],T)+\mathcal{F}_{\text{ext}}([\rho],T)+\mathcal{F}_{\text{exc}}([\rho],T),$ (16)
where $\mathcal{F}_{\text{id}}([\rho],T)=\frac{1}{\beta}\int d^{3}\mathbf{r}\rho(\mathbf{r})(\ln(\Lambda^{3}\rho(\mathbf{r}))-1)$, $\beta=\left(1/{k_{B}T}\right)$ and $\Lambda$ is the thermal wavelength, and $\mathcal{F}_{\text{ext}}([\rho])=\int d^{3}\mathbf{r}\rho(\mathbf{r})V_{\text{ext}}(\mathbf{r})$.
The challenge is to approximate $\mathcal{F}_{\text{exc}}[\rho]$ which represents the inter-particle interactions as a functional of the non-local density. Once such an $\mathcal{F}[\rho]$ is (approximately) obtained, the equilibrium density $\rho_{\text{eq}}(\rho)$ from Eq. 14 can be formulated as the following Euler-Lagrange equation,
$\rho_{\text{eq}}(\mathbf{r})=\frac{1}{\Lambda^{3}}\exp\left(\beta\mu-\left.\beta\frac{\delta\mathcal{F}^{exc}[\rho]}{\delta\rho(\mathbf{r})}\right|_{\rho=\rho_{\text{eq}}}-\beta V_{\text{ext}}(\mathbf{r})\right).$ (17)
This self-consistency relation can be solved by means of a recursive Picard iteration scheme to compute the equilibrium density starting from any arbitrary initial density. It is also worth noting that cDFT can be interpreted as a grand-canonical framework in which the system is in thermal and diffusive equilibrium with a particle bath at chemical potential $\mu$ and temperature $T$.
Figure 5: An overview of single-body direct correlation matching.
A.1 Single-Body Direct Correlation Matching
In addition to pair-correlation matching [10], we can also train a neural free-energy functional ($\mathcal{F}_{\theta,\text{exc}}^{(1)}$) using single-body direct correlation matching. This is done by minimizing the error between the target ${\delta\mathcal{F}_{\text{exc}}}/{\delta\rho(z_{i})}$ and the estimate $(1/{\Delta z}){\partial\mathcal{F}_{\theta,\text{exc}}^{(1)}}/{\partial\rho_{i}}$. This functional derivative of the excess free-energy is the one-body direct correlation function, and computing the target samples requires simulating inhomogeneous equilibrium densities under a range of external potentials.
Appendix B Dynamical Density Functional Theory
We provide here a phenomenological derivation for classical dynamical density functional theory (DDFT) starting from cDFT, originally presented by Evans [3]. DDFT can also alternatively be derived from Langevin equations [11], Smoluchowski equations [12], as well as from a projection operator formalism [39]. Consider the continuity equation for conserved $\rho$,
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}+\nabla\cdot\mathbf{J}(\mathbf{r},t)=0.\$ (18)
Considering the flux $\mathbf{J}(\mathbf{r},t)=\rho(\mathbf{r},t)\mathbf{v}(\mathbf{r},t)$,
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}+\nabla\cdot\left(\rho(\mathbf{r},t)\mathbf{v}(\mathbf{r},t)\right)=0,\$ (19)
where the average velocity $\mathbf{v}(\mathbf{r},t)=-\Gamma\nabla\mu(\mathbf{r},t)$ with the mobility coefficient $\Gamma$. Expanding this we get,
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}-\Gamma\nabla\cdot\left(\rho(\mathbf{r},t)\nabla\mu(\mathbf{r},t)\right)=0.\$ (20)
From cDFT, we know at equilibrium $\mu={\delta\mathcal{F}[\rho(\mathbf{r})]}/{\delta\rho(\mathbf{r})}$. This equilibrium density functional can also be applied to DDFT close to equilibrium. This leads to,
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}=\Gamma\nabla\cdot\left[\rho(\mathbf{r},t)\nabla\left(\frac{\delta\mathcal{F}[\rho(\mathbf{r},t)]}{\delta\rho(\mathbf{r},t)}\right)\right].$ (21)
Re-arranging the terms and expanding the free-energy functional we get,
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}$ $=\Gamma\nabla\cdot\Bigg{[}\rho(\mathbf{r},t)\nabla\Bigg{(}\frac{\delta\mathcal{F}_{\text{exc}}[\rho]}{\delta\rho(\mathbf{r},t)}+V_{\text{ext}}(\mathbf{r},t)$
$\qquad\qquad+\frac{1}{\beta}\ln{\left(\Lambda^{3}\rho(\mathbf{r},t)\right)}\Bigg{)}\Bigg{]}.$ (22)
Appendix C MOF-like planar potentials
Here we describe our procedure for obtaining the $1$D external potentials from metal-organic frameworks (MOF) for the breakthrough simulation experiment in the main paper.
We use the IRMOF-1 and IRMOF-10 structures from the IRMOF series of crystal structures [36], which have the same topology but differing pore sizes. To convert their 3D structures into effective 1D external potentials suitable for DDFT, we employed a simulation-based force sampling approach using LAMMPS [29]. Starting from CIF files of the MOFs, we generated compatible data files for LAMMPS with appropriate force field assignments and topology using lammps-interface [40]. The pore axis of each MOF was aligned along the z-direction, and a single LJ-type guest particle was introduced and displaced incrementally along its pore axis and the total force acting on it was recorded at each step. This produced a raw force profile as a function of z-position which was then converted to reduced LJ units and numerically integrated (in reverse) to generate a dimensionless potential profile. The resulting potentials were clipped and centered to remove artifacts at the boundaries and padded with zeros to enforce a flat, finite domain. The final 1D potentials, expressed as functions interpolated over a grid, serve as external fields $V_{\text{ext}}$ that approximate the confinement environment of the original MOF pores.
Figure 6: IRMOF-1 (top) and IRMOF-10 (bottom) and their corresponding effective 1D potentials computed along the pore axes.
Appendix D Runtime comparisons
The below table is a comparison of approximate runtimes between BD simulation and our DDFT implementation for open and closed systems across two different grid sizes. While each DDFT simulation directly produces smooth density profiles, each BD simulation needs to be ensemble averaged (over 5000 and 500 runs for closed and open resp.), to obtain comparable density profiles. Both simulations were run on a single-core of Intel(R) Xeon(R) W-2123 CPU @ 3.60GHz or comparable CPU. Note that DDFT uses $\Delta t=10^{-3}\tau$ for grid spacing ${\sigma}/{32}$ and $\Delta t=10^{-4}\tau$ for grid spacing ${\sigma}/{100}$, while BD uses $\Delta t=10^{-3}\tau$ for both.
System Grid BD DDFT
Closed (Eq. (M1), ${\sigma}/{32}$ 34 hours 25 secs.
until $t=4\tau$) ${\sigma}/{100}$ 34 hours 5 mins.
Open (Eq. (M2), ${\sigma}/{32}$ 48 hours 40 mins.
until $t=150\tau$) ${\sigma}/{100}$ 48 hours 8 hours
Appendix E Supplemental Movies
We also attach two movie files as supplemental material (also available on the project webpage [41]).
dynamical-density-profiles.mp4: Dynamical version of Fig. (3) showing the density profiles in the main paper.
breakthrough-simulation.mp4: Dynamical version of Fig. (4) showing the breakthrough curves in the main paper.